OpenStudy (anonymous):

I just wanted to check my answer: As more resistors are connected in a parallel circuit, the overall current in the circuit Increases Decreases Stays the same My answer was stay the same

OpenStudy (kmeis002):

dependent on R, and adding R in parallel does change the total R< the total I must change. The question is, do you think the Resistance will increase or descrease?

OpenStudy (anonymous):

I thought they would stay the same depending on the the amount of Ohms that the resister was.

OpenStudy (kmeis002):

Well, the equivalent (total) resistance in a parallel circuit is \[\frac{1}{R_{eq}} = \sum_{n=1}^{k} \frac{1}{R_n} = \frac{1}{R_1}+\frac{1}{R_2}+... \] So, as we add resistance (any finite amount of resistance), the inverse of the total resistance will increase so the total resistance must decrease. The only way the R_eq will remain constant is if we add a resistor with infinite resistance (IE a short or adding none at all).

OpenStudy (anonymous):

Ok I understand now, I was using a simulator for this question, not the formula. When i changed the value of the Ohms in the resistors it did not seem to change the current unless i added other resisters with different Ohms.